OpenStudy (anonymous):

Subtract five and one-fourth minus three and three-eighths. Justify your steps Please help

OpenStudy (♪chibiterasu):

\[5\frac{ 1 }{ 4 } - 3\frac{ 3 }{ 8 } = ?\] We need to make the denominator of 1/4 become 8. We do this by multiplying by 2: 2/2 • 1/4 = 2/8 Now we have: \[5\frac{ 2 }{ 8 } - 3\frac{ 3 }{ 8 } = ?\] Now subtract: 5 - 3 = 2 2 - 3 = -1 The answer would be: \[2\frac{ -1 }{ 8 }\] Which simplifies to -0.25 or 1/4